limx40 sin(4.c) cos(3.c) tan 2x = ? limo–0 2–1 eX – 1 = ? lim.c10+ x ln x : ? S-|(10x4 + 1) dx = ?

Will give thumbs up!
Answer #1

Son Consider given sinux-casix lim xo tanaa lim sinua. Casax a>0 tem 22 sino. Caso temo olo Apply GHRule Ces 48.4 ceas 3x + Sgiven ex_) lm no 2-1 let lim ed-1 el = y 1-1 -1 x=0 2- 0-1 o Aug sejurk clim nyot lim (xjupe let af Xyot lim hao (0+). In co8) given (lox +) dx ། ( ) kel C + 2K + [ ༡༣) = [ 2,++] - [- 4 - ] = 3 - [- ] - 3 + ༦